OpenStudy (christos):

Use the coterminal theorem to evaluate cos420*

OpenStudy (anonymous):

Cos(420) =Cos(420-360)=Cos(60) =1/2

OpenStudy (christos):

Why cos(60) = 1/2 where can I find that information

OpenStudy (anonymous):

on trig table or on calculator

OpenStudy (anonymous):

Best_mathematician please don't just copy my answer...

OpenStudy (anonymous):

oh sry ya he got it first but i didnt copy yours...dont blame and ya i dont want stupid medals give it to him